# Wilsonian Renormalization of Neural Network Gaussian Processes

### v1

@@ Abstract
Separating relevant and irrelevant information is key to any modeling process or scientific inquiry. Theoretical physics offers a powerful tool for achieving this in the form of the renormalization group (RG). Here we demonstrate a practical approach to performing Wilsonian RG in the context of Gaussian Process (GP) Regression. We systematically integrate out the unlearnable modes of the GP kernel, thereby obtaining an RG flow of the Gaussian Process in which the data plays the role of the energy scale. In simple cases, this results in a universal flow of the ridge parameter, which becomes input-dependent in the richer scenario in which non-Gaussianities are included. In addition to being analytically tractable, this approach goes beyond structural analogies between RG and neural networks by providing a natural connection between RG flow and learnable vs. unlearnable modes. Studying such flows may improve our understanding of feature learning in deep neural networks, and identify potential universality classes in these models.
1 Introduction
The renormalization group (RG) is one of the pillars of theoretical physics peskin1995 ; cardy1996scaling . Besides providing an analytical toolkit for tackling large-scale interacting systems (e.g. WILSON1974 ; Gross1973 ; KT1973 ), RG underpins concepts such as universality, irrelevance, and scaling. In complex physical systems, RG is crucial for our ability to model and predict emergent phenomena. It is therefore interesting to ask how may this conceptual framework generalizes to complex virtual systems, in particular deep neural networks (DNNs).
From an experimental point of view, DNNs exhibit various phenomena common to RG domains. They are large-scale systems kaplan2020scaling , exhibiting collective phenomena where the importance of each specific parameter is minute molchanov2017pruning . In realistic settings, DNNs often show robust scaling laws in performance versus dataset size or the number of tunable parameters bahri2021explaining ; kaplan2020scaling . In a related manner, data-data covariance matrices, associated with networks’ outputs, also exhibit power law decaying spectra bahri2021explaining ; bordelon2021spectrum similar in spirit to Zipf’s law. Finally, universality seems present to some degree, as often large microscopic/mechanistic changes, such as a change of activation function or loss function, result in minor performance changes Novak2018BayesianDC ; Lee2020FiniteVI .
Here we suggest a distinct route where RG may come into quantitative use. To this end, we model DNNs as Gaussian Processes (GPs), a description which becomes exact in a suitable infinite-overparametrization limit Neal ; jacot2018neural ; lee2018deep ; matthews2018gaussian ; Cohen_2021 and applies to a wide range of DNNs. It is also extendable to finite DNNs via adaptive kernel approaches LiSompolinsky2021 ; Seroussi2023 ; ariosto2022statistical . In the GP viewpoint, the action governing the theory is quadratic and described by a kernel which encodes the network’s architecture. Diagonalizing this kernel, one obtains a set of eigenfunctions. Fluctuations of network outputs along eigenfunctions with low kernel eigenvalues are highly suppressed, and are thus analogous to “high-energy” modes in RG. [^1] To avoid confusion, note that the language “high” and “low energy” stems from the field-theoretic practice of working in frequency space; here, the kernel eigenmodes are akin to wavelengths, i.e., “high-energy” modes mean modes with low eigenvalues. Such modes, if present in the target, are also harder to learn Rasmussen ; Cohen_2021 ; canatar . Notably, while the GP theory is quadratic and, in this sense, exactly solvable, it contains randomness (or quenched disorder) associated with the specific choice of training set. Averaging over the latter leads to a non-quadratic/interacting replicated field theory Cohen_2021 ; canatar .
Our contributions.
@@ 1 Introduction
In this work, we apply Wilsonian RG to Gaussian Processes regression by iteratively integrating out these high-energy GP modes. Sec. 2 presents general layout for this approach. We show that in various relevant scenarios, such a procedure is tractable and results in a renormalized GP theory. When the high-energy eigenfunctions are uncorrelated, in a manner to be made precise below, this amounts to a scalar renormalization of the loss or equivalently the ridge parameter; this is developed in sec. 3. When eigenfunction correlations are present, a further input-dependent/spatial weighting of the loss appears as well as potential shifts to the target; this is demonstrated in sec. 4. The resulting RG flow proceeds until it reaches the first learnable eigenfunction. Thus, our lower RG scale (i.e., the infrared) naturally depends on the type and amount of data. In cases where high-energy modes have sizable collective fluctuations, the RG flow ends in a weakly interacting regime amenable to perturbative methods Cohen_2021 . Alternatively, it may end in a strongly interacting theory which, in generic high dimensional scenarios, is amenable to saddle point techniques canatar . Lastly, we point out that for power-law decaying kernel spectra, often found in real-world data, we obtain an RG flow towards the weakly interacting regime. By relying on rather general formalism and introducing a precise and practically motivated notion of RG, we hope to provide a stepping stone towards establishing notions of universality in deep learning.
Related works.
Several works have considered connections and analogies between deep learning and RG. For example, Halverson:2020trp ; Erbin_2022 ; erbin2022renormalization ; erbin2023functional examined RG flows in the couplings of a corresponding field theory due to finite-width effects (see also Grosvenor:2021eol ). Similar effects arising layer-by-layer have been discussed in Roberts_2022 from an effective theory perspective. The structural analogy between RG and DNNs was further explored in Erdmenger:2021sot , which tracked the flow of information as a function of depth. These approaches are however primarily concerned with untrained networks and do not provide a relation between RG and learning dynamics via the data. Another line of work Cotler_2023 ; cotler2023renormalizing explored RG-inspired training schemes for diffusion models and optimal transport in DNNs, where updates of network outputs are modeled after Polchinski’s so-called Exact RG (ERG) flow. In addition, berman2022dynamics ; berman2022inverse ; Berman_2023 ; berman2024ncoder construct an information theoretic ERG flow to describe the inverse of Bayesian inference in DNN ensembles. What distinguishes our approach is the natural connection between the energy scale in Wilsonian RG and the data, which allows us to perform RG on the network’s degrees of freedom for networks trained in a standard manner, using the framework in Cohen_2021 .
Finally, as this manuscript was completed, a related work atanasov2024scaling appeared where an approach based on free probability was used to derive “renormalization” of the ridge parameter and other quantities relevant for GP regression with Gaussian features. There, the change to the ridge parameter arises from a self-energy-type diagrammatic resummation; however, while diagrammatic resummations are of course part of the toolkit of field theory, central RG notions such as a cut-off or a flow of couplings were not made explicit. The current manuscript goes beyond the case of Gaussian features, and develops a direct Wilsonian RG framework which further generalizes to finite-width effects systematically.
2 Equivalent Kernel Limit of the Replica Action
@@
To provide a relatively self-contained derivation, we first define Gaussian Process Regression, its relations to neural networks, and replica techniques used to obtain the average predictions over ensembles of datasets. In addition to being a well-established tool in machine learning, Gaussian Process Regression describes both the lazy learning or kernel regime of an infinite-width neural network lee2018deep ; jacot2018neural ; Cohen_2021 , alongside extensions to finite networks via the aforementioned adaptive kernel approaches.
@@
The output of such random networks behaves as a Gaussian Process with zero mean and a covariance function $K(x,x^{\prime})$, called the kernel. Here, the training data, $x,x^{\prime}\in\mathbb{R}^{d}$, are drawn from a data measure [^2] By this we simply mean a measure over the data that admits an interpretation as a probability distribution. That is, let $\mathrm{d}\mu_{x}$ be a measure on $x\in\mathbb{R}^{d}$ such that for some probability distribution $p_{\mathrm{data}}(x)$, we have $\int\!f(x)\mathrm{d}\mu_{x}=\int\!f(x)p_{\mathrm{data}}(x)\mathrm{d}x$ for $f:\mathbb{R}^{d}\to\mathbb{R}$. In a slight abuse of notations, we will refer to both the measure $\mathrm{d}\mu_{x}$ and the distribution $p_{\mathrm{data}}(x)$ as the data measure. $p_{\mathrm{data}}(x)$; the total training set consists of $n$ such draws, and is denoted by the $n\times d$ matrix $X_{n}$. We further consider a regression problem with a target, $y(x)$. A standard result in GP regression states that Rasmussen the average predictor on a test point, $\tilde{f}(x_{*})$, given $X_{n}$ is
@@
Since the average predictor (1) is conditional on $X_{n}\!\sim\!p_{\mathrm{data}}(x)$, it is sensitive to the details of the specific dataset we drew. To remove this dependence, we consider the dataset-averaged predictor by computing the expectation value with respect to $p_{\mathrm{data}}(x)$. The most general way of doing so in a tractable manner uses a “grand-canonical” trick Malzahn ; Cohen_2021 , which assumes that the number of training points is drawn from a Poisson distribution with expectation value $\eta$,
@@
such that $\eta$ now controls the average number of data points. Notably, due to the nature of the Poisson distribution, in the relevant limit of large $\eta$, the fluctuations become increasingly negligible compared to the average.
@@
This replica approach was first applied to GP regression in Malzahn , and was used in Cohen_2021 to obtain
@@
$S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})-\eta\int\!\mathrm{d}\mu_{x}\,e^{-\sum_{m=1}^{M}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2\sigma^{2}}}\leavevmode\nobreak\ ,$ (5)
@@
Although our initial field theory was Gaussian, averaging over the datasets resulted in a non-quadratic action, due to “quenched-disorder" induced by the training set. A simple yet useful limit Sollich2004 of the above theory, known as the Equivalent Kernel limit, is obtained when $\eta,\sigma^{2}\rightarrow\infty$ with $\eta/\sigma^{2}$ fixed. In this case the exponential on the right-hand side of (5) can be Taylor expanded, so that, at leading order, we obtain a non-interacting theory,
$\eta+S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$ (6)
@@
see Appendix A. Note that the path integral measure also decouples into a measure over the feature modes, $\mathcal{D}f_{m}\sim\prod_{k}\mathcal{D}f_{mk}$, up to a constant Jacobian factor. Given this action, one can then show that the average prediction is Cohen_2021 ; Rasmussen
@@
3 Renormalizing Gaussian features
Formally, our Wilsonian renormalization procedure can be explained very simply: we wish to integrate out all modes above some cutoff $\kappa$, in order to obtain an effective action in terms of only the modes with $k<\kappa$. Accordingly, we split the spectral decomposition (9) into
$f_{m<}(x)\coloneqq\sum_{k\leq\kappa}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,\qquad\quad f_{m>}(x)\coloneqq\sum_{k>\kappa}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,$ (13)
@@ 3 Renormalizing Gaussian features
$y_{<}(x)\coloneqq\sum_{k\leq\kappa}y_{k}\phi_{k}(x)\leavevmode\nobreak\ ,\qquad\quad y_{>}(x)\coloneqq\sum_{k>\kappa}y_{k}\phi_{k}(x)\leavevmode\nobreak\ .$ (14)
@@
$-\eta\int\!\mathrm{d}\mu_{x}\,\exp\!\left(-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\left(f_{m<}-y_{<}\right)+\left(f_{m>}-y_{>}\right)\right]^{2}\right)\leavevmode\nobreak\ ,$
@@
$\langle Z^{M}\rangle_{\eta}$ $=e^{-\eta}\int\!\prod_{m}\mathcal{D}f_{m<}\,e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}$ (16)
@@
$S_{\mathrm{eff}}[f_{m<}]\coloneqq-\ln\left(e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}\right)\leavevmode\nobreak\ ,$ (17)
with $\mathcal{D}f_{m<}\coloneqq\prod_{k\leq\kappa}\mathrm{d}f_{mk}$, and similarly for $\mathcal{D}f_{m>}$. In field-theoretical language, the effective action corresponds to the low-energy (cf. footnote 1) effective action that describes the system up to the scale set by the cutoff, $\kappa$. From a Bayesian perspective, this amounts to marginalizing over the hidden degrees of freedom that are inaccessible to the observer.
@@
$1=\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ ,$ (18)
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\underbrace{\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)}_{P[\varphi_{k}]}\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]^{2}\right\}$ (19)
$=-\eta\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,P[\varphi_{k}]\,\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\varphi_{k}\right]^{2}\right\}\leavevmode\nobreak\ ,$
where we have introduced the feature space distribution
$P[\varphi_{k}]\coloneqq\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}x\,p_{\mathrm{data}}(x)\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ .$ (20)
@@
$\Phi_{m<}\coloneqq\sum_{k\leq\kappa}\left(f_{mk}-y_{k}\right)\varphi_{k}\leavevmode\nobreak\ ,\qquad\quad\Phi_{m>}\coloneqq\sum_{k>\kappa}\left(f_{mk}-y_{k}\right)\varphi_{k}\leavevmode\nobreak\ .$ (21)
@@
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak$
$=-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak\ ,$
@@
We can now proceed via a two-stage process, wherein we first integrate out the greater feature modes with the GP modes held fixed, and then integrate out the greater GP modes ($f_{mk}$ with $k>\kappa$). To this end, we approximate $P[\varphi]$ as a multivariate Gaussian distribution. We note in passing that a weaker requirement, namely $\Phi_{>}$ is Gaussian under the $p_{\mathrm{data}}(x)$ data measure, also suffices; cf. canatar . At a technical level, this assumption allows us to perform the spatial integration in $S_{\mathrm{int}}$ and recast the action in terms of $f_{mk}$ alone.

### v2

@@ Abstract
Separating relevant and irrelevant information is key to any modeling process or scientific inquiry. Theoretical physics offers a powerful tool for achieving this in the form of the renormalization group (RG). Here we demonstrate a practical approach to performing Wilsonian RG in the context of Gaussian Process (GP) Regression. We systematically integrate out the unlearnable modes of the GP kernel, thereby obtaining an RG flow of the GP in which the data sets the IR scale. In simple cases, this results in a universal flow of the ridge parameter, which becomes input-dependent in the richer scenario in which non-Gaussianities are included. In addition to being analytically tractable, this approach goes beyond structural analogies between RG and neural networks by providing a natural connection between RG flow and learnable vs. unlearnable modes. Studying such flows may improve our understanding of feature learning in deep neural networks, and enable us to identify potential universality classes in these models.
preprint: arXiv:2405.06008
I Introduction
Deep neural networks (DNNs) are showing an increasing amount of universality across tasks [1], dataset sizes, and computational power [2]. In simple asymptotic limits, power-law scaling of DNN performance [2] can be traced to Zipf’s law behaviour on the underlying datasets [3, 4]. In other, more non-linear limits, the effects of non-linearities can often be absorbed as a change of a single scale parameter [5, 6, 7]. As DNNs are large-scale systems in which the importance of each specific parameter is minute [8], some self-averaging effects leading to universality are perhaps expected, yet their extent remains to be explored both theoretically and empirically.
In physics, scale-free behavior, renormalization of parameters, and universality, are key aspects of the Renormalization Group (RG), one of the pillars of theoretical physics [9, 10]. At its core, RG is an analytical technique that generates a continuous transformation (flow) between models/physical systems. This transformation is defined such that all models along a given flow agree on low-resolution/coarse-grained observables. Critical theories, showing many of the aforementioned scale-free phenomena, are described by points on this flow at which the correlation length diverges. Besides providing an analytical toolkit for tackling large-scale interacting systems (e.g., [11, 12, 13]), RG is an inherent part of modelling systems (via notions of relevancy), and offers qualitative insights into the role of non-linearities via the regularities of low-dimensional flows.
In the context of increasing efforts to make AI safer and more predictable, the aforementioned scale-free universal behaviors of DNNs are encouraging and invite interpretations in terms of some renormalization group flow. At a practical level, there are various different procedures for implementing RG, the choice of which is best determined by physical criteria, chiefly: (i) What is the scale? This determines which observables are being integrated or coarse-grained out, and which remain. (ii) What is the UV theory on which one is applying RG? (iii) Is the induced RG flow tractable, or does it make some quantities more tractable?
An early line of work sought to augment (or interpret) the action of a layer in deep neural networks so that it coarse-grains the data distribution, thus drawing an analogy with RG [14, 15, 16]. Similarly, the action of a diffusion model can be interpreted as inverting an RG flow on the data distribution [17]. The UV theories here are thus the data distribution and the IR theory consists of the relevant features in the data for the layer action and white noise data for the diffusion action. Another line of work [18, 19] views the distribution on functions induced by a random DNN as the UV theory, and uses RG formulations to treat non-Gaussian (finite-width) effects [20, 21, 22] in this distribution. However, the notion of scale and the relation to actual DNN priors remains quite varied across these works. Finally, another set of works relating RG flows and Bayesian Inference view both as forms of augmenting a probability distribution [23, 24, 25, 26, 27]. In this approach however, reducing data generates the RG flow, so the true data-generated model (or target function) is the UV theory, while the IR theory is a random DNN, so that observables that are consistently reproduced along the flow are the most unlearnable modes. To the best of our understanding, these works have yet to analytically predict the generalization performance of neural networks.
Here we formulate a distinct route where RG may come into quantitative use in predicting DNN performance. Our UV theory is the distribution over outputs generated by an ensemble of trained neural networks at finite amount of data, while the IR is the theory for the learnable modes. To generate an RG flow, we integrate out unlearnable output modes and track their renormalizing effect on the learnable modes. To illustrate our approach, we focus on a suitable infinite-overparameterization limit [28, 29, 30, 31, 32] (the Gaussian Process limit) which applies to a wide range of DNNs. We focus on using RG to treat the non-linearities induced by data-averaging/finite-sample-size effects rather than those induced by finite-width, though our formalism naturally extends to that case as well.
@@ I Introduction
We show that in various relevant scenarios, our RG procedure is tractable and results in a renormalized Gaussian Process (GP) theory. When the unlearnable modes are uncorrelated, in a manner to be made precise below, this amounts to a scalar renormalization of the loss or equivalently the ridge parameter, thus providing an RG interpretation for the results of [5]; this is developed in sec. III. When correlations between unlearnable modes are present, a further input-dependent/spatial weighting of the loss appears as well as potential shifts to the target; this is demonstrated in sec. IV. The resulting RG flow proceeds until it reaches the first learnable output modes. Thus, our lower RG scale (i.e., the infrared) naturally depends on the type and amount of data. In cases where unlearnable modes have sizable collective fluctuations, the RG flow ends in a weakly interacting regime amenable to perturbative methods [32]. Alternatively, it may end in a strongly interacting theory which, in generic high dimensional scenarios, is amenable to saddle point techniques [5]. Lastly, we point out that for power-law decaying kernel spectra, often found in real-world data, we recover scaling-law results [3], thereby establishing a concrete link between RG and scaling-laws.
By relying on rather general formalism and introducing a down-to-earth practically motivated notion of RG, we hope to provide a stepping stone towards establishing notions of universality in deep learning.
II Equivalent Kernel Limit of the Replica Action
@@
To provide a relatively self-contained derivation, we first define Gaussian Process Regression, its relations to neural networks, and replica techniques used to obtain the average predictions over ensembles of datasets. In addition to being a well-established tool in machine learning, Gaussian Process Regression describes both the lazy learning or kernel regime of an infinite-width neural network [30, 29, 32], alongside extensions to finite networks via approaches detailed in [6, 33, 34].
@@
The output of such random networks behaves as a Gaussian Process with zero mean and a covariance function $K(x,x^{\prime})$, called the kernel. Here, the training data, $x,x^{\prime}\in\mathbb{R}^{d}$, are drawn from a data measure [^1] By this we simply mean a measure over the data that admits an interpretation as a probability distribution. That is, let $\mathrm{d}\mu_{x}$ be a measure on $x\in\mathbb{R}^{d}$ such that for some probability distribution $p_{\mathrm{data}}(x)$, we have $\int\!f(x)\mathrm{d}\mu_{x}=\int\!f(x)p_{\mathrm{data}}(x)\mathrm{d}x$ for $f:\mathbb{R}^{d}\to\mathbb{R}$. In a slight abuse of notations, we will refer to both the measure $\mathrm{d}\mu_{x}$ and the distribution $p_{\mathrm{data}}(x)$ as the data measure. $p_{\mathrm{data}}(x)$; the total training set consists of $n$ such draws, and is denoted by the $n\times d$ matrix $X_{n}$. We further consider a regression problem with a target, $y(x)$. A standard result in GP regression states that [35] the average predictor on a test point, $\tilde{f}(x_{*})$, given $X_{n}$ is
@@
Since the average predictor (1) is conditional on $X_{n}\!\sim\!p_{\mathrm{data}}(x)$, it is sensitive to the details of the specific dataset we drew. To remove this dependence, we consider the dataset-averaged predictor by computing the expectation value with respect to $p_{\mathrm{data}}(x)$. The most general way of doing so in a tractable manner uses a “grand-canonical” trick [36, 32], which assumes that the number of training points is drawn from a Poisson distribution with expectation value $\eta$,
@@
such that $\eta$ is now the average number of data points. Notably, due to the nature of the Poisson distribution, in the relevant limit of large $\eta$, the fluctuations become increasingly negligible compared to the average.
@@
This replica approach was first applied to GP regression in [36], and was used in [32] to obtain
@@
$S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (5)
$\qquad-\eta\int\!\mathrm{d}\mu_{x}\,e^{-\sum_{m=1}^{M}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2\sigma^{2}}}\leavevmode\nobreak\ ,$
@@
Although our initial field theory was Gaussian, averaging over the datasets resulted in a non-quadratic action, due to “quenched-disorder” induced by the training set. A simple yet useful limit [37] of the above theory, known as the Equivalent Kernel (EK) limit, is obtained when $\eta,\sigma^{2}\rightarrow\infty$ with $\eta/\sigma^{2}$ fixed. In this case the exponential on the right-hand side of (5) can be Taylor expanded, so that, at leading order, we obtain a non-interacting theory,
$\eta+S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (6)
$\qquad+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$
@@
see Appendix A. Note that the path integral measure also decouples into a measure over the feature modes, $\mathcal{D}f_{m}\sim\prod_{k}\mathcal{D}f_{mk}$, up to a constant Jacobian factor. Given this action, one can then show that the average prediction is [32, 35]
@@
III Renormalizing Gaussian features
Formally, our Wilsonian renormalization procedure can be explained very simply: we wish to integrate out all modes above some cutoff $\kappa$, in order to obtain an effective action in terms of only the modes with $k\leq\kappa$. Accordingly, we split the spectral decomposition (9) into
$f_{m<}(x)\coloneqq\sum_{k\leq\kappa}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,\quad f_{m>}(x)\coloneqq\sum_{k>\kappa}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,$ (13)
@@ III Renormalizing Gaussian features
$y_{<}(x)\coloneqq\sum_{k\leq\kappa}y_{k}\phi_{k}(x)\leavevmode\nobreak\ ,\quad y_{>}(x)\coloneqq\sum_{k>\kappa}y_{k}\phi_{k}(x)\leavevmode\nobreak\ .$ (14)
@@
$-\eta\int\!\mathrm{d}\mu_{x}\,\exp\!\Bigg{(}-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\left(f_{m<}-y_{<}\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+\left(f_{m>}-y_{>}\right)\right]^{2}\Bigg{)}\leavevmode\nobreak\ ,$
@@
$\langle Z^{M}\rangle_{\eta}$ $=e^{-\eta}\int\!\prod_{m}\mathcal{D}f_{m<}\,e^{-S_{0}[f_{m<}]}$ (16)
×$\qquad\times\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}$
@@
$S_{\mathrm{eff}}[f_{m<}]\coloneqq$ $-\ln\Bigg{(}e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,$ (17)
×$\qquad\qquad\times e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}\Bigg{)}\leavevmode\nobreak\ ,\leavevmode\nobreak$
with $\mathcal{D}f_{m<}\coloneqq\prod_{k\leq\kappa}\mathrm{d}f_{mk}$, and similarly for $\mathcal{D}f_{m>}$. In field-theoretical language, the effective action corresponds to the low-energy [^4] To avoid confusion, note that the language “high” and “low energy” stems from the field-theoretic practice of working in frequency space; here, the kernel eigenmodes are akin to wavelengths, i.e., “high-energy” modes mean modes with low eigenvalues. effective action that describes the system up to the scale set by the cutoff, $\kappa$. From a Bayesian perspective, this amounts to marginalizing over the hidden degrees of freedom that are inaccessible to the observer.
@@
11 11 $=\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ ,$ (18)
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\mathrm{d}\mu_{x}\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)$ (19)
×$\qquad\qquad\times\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]^{2}\right\}$
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi]\,$
×$\qquad\qquad\times\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\varphi_{k}\right]^{2}\right\}\leavevmode\nobreak\ ,$
where in the second line we implicitly defined the joint distribution over all $\varphi_{k}$’s as well as the new product measure, respectively,
$P[\varphi]\leavevmode\nobreak\ \coloneqq$ $P[\varphi_{1},\varphi_{2},...]=\int\!\mathrm{d}\mu_{x}\,\prod^{\infty}_{k=1}\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ ,\qquad$ (20)
$\mathcal{D}\varphi\leavevmode\nobreak\ \coloneqq$ $\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\leavevmode\nobreak\ .$ (21)
@@
The features $\varphi$ can be divided into two sets $\varphi_{>}$ and $\varphi_{<}$ above and below the cutoff $\kappa$, respectively. Therefore, the joint distribution $P[\varphi]$ can be divided into probability distributions over these two sets via the conditional relation,
$P[\varphi]=P[\varphi_{<},\varphi_{>}]=P[\varphi_{<}]P[\varphi_{>}|\varphi_{<}]\leavevmode\nobreak\ .$ (22)
@@
$\Phi_{m<}\coloneqq\sum_{k\leq\kappa}\left(f_{mk}-y_{k}\right)\varphi_{k}\leavevmode\nobreak\ ,\quad\Phi_{m>}\coloneqq\sum_{k>\kappa}\left(f_{mk}-y_{k}\right)\varphi_{k}\leavevmode\nobreak\ .$ (23)
@@
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,\exp\!\Bigg{\{}-\frac{1}{2\sigma^{2}}\big{(}\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2\Phi_{<}^{\top}\Phi_{>}\big{)}\Bigg{\}}\leavevmode\nobreak$
$=-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$
×$\times\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak\ ,$
@@
We can now proceed via a two-stage process, wherein we first integrate out the greater feature modes with the GP modes held fixed, and then integrate out the greater GP modes ($f_{mk}$ with $k>\kappa$). To this end, we approximate $P[\varphi]$ as a multivariate Gaussian distribution. We note in passing that a weaker requirement, namely $\Phi_{>}$ is Gaussian under the $p_{\mathrm{data}}(x)$ data measure, also suffices; cf. [5]. At a technical level, this assumption allows us to perform the spatial integration in $S_{\mathrm{int}}$ and recast the action in terms of $f_{mk}$ alone.